OpenStudy (anonymous):

Which phrase best describes the location of 13/8 on a number line? @StudyGuidance

OpenStudy (anonymous):

ok, what number line?

OpenStudy (anonymous):

are you lagging too?

OpenStudy (anonymous):

@trinajoy

OpenStudy (anonymous):

Ya and there is no numberline

OpenStudy (anonymous):

ok gimme a sec

OpenStudy (anonymous):

1 and 5/8

OpenStudy (anonymous):

1 5/8 should be your awnser

OpenStudy (anonymous):

be right back

OpenStudy (anonymous):

There is no 1 5.8 these are the answers

OpenStudy (texaschic101):

there is no 1 and 5/8 ? .....what about 1.625 ??

OpenStudy (anonymous):

1 5/8 is the same as 1.625

OpenStudy (texaschic101):

I know...I just thought since there was no 1 5/8, it might be in decimal form

OpenStudy (anonymous):

oh ok well what are the awnser choices then @trinajoy

OpenStudy (texaschic101):

yes...what are they...that would really help to know that

OpenStudy (anonymous):

@\[A.~At~~\frac{ 3 }{ 8}{ 1 } ~~~B.\frac{ 1 }{ 2 }{ 1 }\]

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

sorry both of them say at

OpenStudy (anonymous):

C. closer to 1 then 2 D.closer to 2 then 1.

OpenStudy (texaschic101):

closer to 2 then 1...because 13/8 = 1 5/8 = 1.625...which is closer to 2 then 1

OpenStudy (anonymous):

ya same

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

ive been on this website for less then 10 min now and ive already been fanned and medaled and asked for help more then 15 times

OpenStudy (anonymous):

yes D

OpenStudy (anonymous):

I have another question

OpenStudy (texaschic101):

ask away

OpenStudy (texaschic101):

@StudyGuidance ...you will reach 99 before you know it

OpenStudy (anonymous):

Thank you for the support :)

OpenStudy (anonymous):

\[Which~ improper ~fraction ~is ~equal~ ~the~mixed~number \frac{ 3 }{ 8 }{ 3 } ? \]

OpenStudy (anonymous):

texas actualy before i had an account i used to look around for awnsers to my questions and you have awnserd more then half of them @texaschic101

OpenStudy (anonymous):

it is 27/8

OpenStudy (anonymous):

1 sec someine needs me

OpenStudy (texaschic101):

mixed number to improper fraction 3 3/8 -- take the whole number (3) and multiply it by the denominator (8), which equals 24, then add that to the numerator (3), which equals 27...now put that over the original denominator (8) giving you : 27/8

OpenStudy (anonymous):

\[A.\frac{ 14 }{ 8 }~~~~~B.\frac{ 17 }{ 8 }~~~~~C.\frac{ 27 }{ 8 }~~~~~D.\frac{ 33 }{ 8 }\]

OpenStudy (texaschic101):

It is C

OpenStudy (anonymous):

Thank you :)

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

i took this exact same test so yeah the is 100% C

OpenStudy (anonymous):

i have 60 people ask for help in 5 min i was off :3